OpenStudy (anonymous):

x^2-9/x^2+2x * x+2/x-3 @kc_kennylau

OpenStudy (anonymous):

Alot tougher huh :)

OpenStudy (kc_kennylau):

\[\Large\frac{x^2-9}{x^2+2x}\cdot\frac{x+2}{x-3}\] Why can't you just put parentheses lol

OpenStudy (kc_kennylau):

ok factorize \(x^2-9\) first by using \(a^2-b^2=(a-b)(a+b)\) :)

OpenStudy (kc_kennylau):

Then factorize \(x^2+2x\) by taking out the common factor :)

OpenStudy (anonymous):

Oh the perfect square formula

OpenStudy (kc_kennylau):

oh you call it the perfect square formula? interesting name :p

OpenStudy (anonymous):

(x+3)(x-3)

OpenStudy (anonymous):

right

OpenStudy (kc_kennylau):

then factorize \(x^2+2x\) by taking out the common factor :)

OpenStudy (anonymous):

x(x+2)

OpenStudy (anonymous):

ahhhh

OpenStudy (anonymous):

now they cancel!

OpenStudy (kc_kennylau):

yep :D So now we got \(\dfrac{(x+3)(x-3)}{x(x+2)}\cdot\dfrac{x+2}{x-3}\)

OpenStudy (kc_kennylau):

yep they cancel :D

OpenStudy (anonymous):

AWESOME!

OpenStudy (anonymous):

i love math!

OpenStudy (kc_kennylau):

good :D keep up this passion :D

OpenStudy (anonymous):

alirghty! college algebra final here i come!

OpenStudy (anonymous):

Ill tag you if i need any more help.

OpenStudy (kc_kennylau):

ok :D